I have a problem on my study guide that i'm having trouble. The question is," M(5,3) is the midpoint of RS. If S has coordinates (9,12), find the coordinates of R." I know how to solve a midpoint problem I just don't know how to write this equation.
Answer by richwmiller(17219)

Just do the reverse.
(X1+X2)/2=Xm
(Y1+Y2)/2=Ym
5,3
9,12
9+X2/2=5
9+X2=1
12+Y2/2=3
12+Y2=6
Y=-6
endpoint =(1,-6)
check with line equation
y = (9 x)/4-33/4
-6 = (9 *1)/4-33/4
-24=9-33